# A Joint Representation Using Continuous and Discrete Features for Cardiovascular Diseases Risk Prediction on Chest CT Scans

### v1

@@ 0.6 Deep continuous feature and discrete biomarker extraction
The extraction of deep continuous features relies on a pre-trained CVD risk prediction model, which has been trained on the LDCT-NLST training set. The model is based on a progressive coarse-to-fine framework. In the coarse stage (heart localization), a lightweight cardiac segmentation model is applied to localize the volume of the region of interest in the heart. The input chest LDCT scan is cropped into a smaller volume to reduce the influence of irrelevant content noise and conserve computational resources. Moving to the fine stage, the cropped volume is used to train a classifier based on ResNet34 [he2016deep], tasked with distinguishing between CVD-Positive and CVD-Negative cases. Ultimately, the outputs from the layer preceding the final fully connected layer of ResNet34 are extracted. These high-dimensional embeddings are rich in features relevant to CVDs, denoted as (, here ), and serve as continuous deep features of our proposed method.
@@
The CT quantitative biomarkers extraction process utilizes four specialized body part segmentation models, each trained on the 400 internal chest CT scans. These models are designed to segment different anatomical structures, including the heart chambers and pericardium, aortic and coronary calcium, aortic structure, and left and right lungs. The foundational architecture for these models is MedFormer [gao2022data]. These four segmentation models are applied to segment the whole LDCT-NLST dataset, with the results being reviewed by two radiologists and revised if necessary. The overall failure rate is less than 0.5. Fig. 4 visually showcases the results obtained from these fully automated body part segmentation models.
@@
We have established stable quantitative measures for each tissue composition based on the automated segmentation results without additional learning or adjustment. A total of quantitative biomarkers are calculated, with certain biomarkers such as the coronary artery calcium score (CACS) and the cardiothoracic ratio (CRT) having established associations with CVD and mortality in previous research [pickhardt2020automated, iacobellis2022epicardial, girardi2021aortic, hemingway1998cardiothoracic, dey2012epicardial]. Others are characterized by physicians using the results of the four segmentation models. All of these biomarkers are scalar and are symbolized as (where ).
@@
(i) Based on the segmentation results of the pericardium, three quantitative biomarkers related to pericardial fat are calculated: PFATV, PFATM, and PFATSTD. PFATV primarily serves to quantify the volume of pericardial fat. Initially, we used the result of pericardial segmentation (here marked in red in 4f) to identify the location of the pericardium on the CT scan. Subsequently, we further delineate the regions of pericardial fat (as depicted in 4b) within the Hounsfield Unit (HU) range of [-190HU, -30HU] [dey2012epicardial]. Following this, we count the number of voxels within the 3D pericardial fat mask (illustrated in 4c), and by multiplying this count with the voxel volume ( resolution resolution resolution), we derive the value for PFATV, expressed in . Finally, we compute the mean and standard deviation of the attenuation intensity values for all voxels within the pericardial fat regions, denoted as PFATM and PFATSTD, respectively.
@@
(ii) Based on the segmentation results of calcifications (Coronary Artery Calcification and Thoracic Aorta Calcification), four calcification-related quantitative biomarkers are calculated: CACS, CACV, ACS, and ACV. CACS and ACS are primarily derived from an area perspective, calculated according to the Agatston Score [agatston1990quantification] for coronary and thoracic aortic calcification scores, respectively. CACV and ACV, on the other hand, are calculated from a volumetric perspective, quantifying the volume of coronary calcification and thoracic aortic calcification, respectively. First, the calcification segmentation results are refined, with only regions with CT attenuation intensity greater than or equal to 130HU retained as the final calcification segmentation mask. Then, calculating calcification scores, the results are uniformly reconstructed to 3mm before computing CACS and ACS to minimize the impact of different CT slice thicknesses. In the calculation of calcification volume, the number of voxels within the 3D coronary calcification mask (green areas in Fig. 4d and Fig. 4e) and thoracic aorta calcification mask (red regions in Fig. 4d and Fig. 4e) are directly tallied, and then multiplied by the voxel volume ( resolution resolution resolution) to obtain CACV and ACV (in ).
@@
(iii) Based on the segmentation results of the thoracic aorta, three quantitative biomarkers characterizing the morphology of the thoracic aorta are computed: ATI, AMD, and AMDSTD. ATI represents the curvature of the thoracic aorta [girardi2021aortic], while AMD and AMDSTD describe the overall maximum diameter of the aorta and its variation along the entire length. To calculate ATI, we first extract the centerline of the thoracic aorta using the segmentation outcome, and then determine the locations of the root point and end point (referred to as Point and Point in Fig. 4f). We then calculate ATI as the ratio of the length of the centerline (in ) to the straight linear distance between the point and the point (also in ). Subsequently, we generate cross-sectional views at 1 intervals along the centerline, with Fig. 4i and Fig. 4j exemplifying the cross-sectional views generated at positions and in Fig. 4(f). For each cross-sectional view, we measure the maximum diameter of the thoracic aorta, denoted as (where = centerline length/interval (1 ) + 1, it represents the total number of cross-sectional views). Finally, we compute AMD as the maximum value in (AMD = ) and ANDSTD as the standard deviation of the diameters in (AMDSTD = ).
@@
(iv) Based on the results of the heart segmentation (cardiac chambers and pericardium), four quantitative biomarkers are computed to characterize the morphology and structure of the heart: CHR, CLD, CSD, and CTR [troxler2018role]. CHR is defined as the ratio of the volume of the cardiac chambers (the green regions in Fig. 4f) to the total volume of the entire heart (the combined red and green areas in Fig. 4f). Using the segmentation results of the heart, we select the maximal four-chambers axial view of the heart and perform an elliptical fitting on this segmented image to derive the cardiac long diameter (CLD) and short diameter (CSD). In Fig. 4f, half of line represents the CSD while half of line represents the CLD. Based on the calculation method described in the previous study [girardi2021aortic], we determine the width of the heart, represented by line in Fig. 4g. Similarly, utilizing the segmentation results of the lungs (as seen in Fig. 4l), we obtain the width of the entire lung, denoted by line in Fig. 4g. Finally, CTR is calculated as the ratio of line to line .
@@
Recent studies have shown that methods that integrate features from various sources into one representation can offer complementary information to each other and lead to better performance [kiela2014learning]. However, not all features contribute to the prediction of the target, and certain artificially designed image quantitative biomarkers may actually hinder performance. Furthermore, the relevance and specific contributions of the features to the output target are typically unknown. Therefore, we have developed a features joint representation module (Fig. 5a) which can help identify which features are most significant for the prediction problem, reinforcing the most relevant features for CVD risk prediction, and suppressing any unnecessary features that could have a negative impact on performance. This module comprises an instance-wise feature-gated mechanism and a soft instance-wise feature interaction mechanism.
@@
To improve the expressive power of the model and better capture the relationship between features and the output target, we apply non-linear processing to each instance-wise feature embedding . However, determining the extent of required non-linear processing remains a complex task. Some studies [pedro2000unified, hawkins2004problem] suggest that simpler models may benefit in datasets with noise. Given that our discrete quantitative biomarkers are automatically derived from the four pre-trained body part segmentation models without any additional adjustment, it implies that the input contains unavoidable noise. In consideration of this, we employ the GRN proposed in [lim2021temporal], which is a notably flexible and simple architecture (Fig. 5b), as the non-linear operation.
Where is the fully-connected Layer, is the Exponential Linear Unit activation function [clevert2015fast], is the Gated Linear Units [dauphin2017language] and is the standard layer normalization [ba2016layer]. During training, dropout is applied, and the dropout rate is set as 0.5. For each instance-wise feature embedding , a non-linear operation is employed by its own GRN:
@@
Once the continuous and discrete features are aligned into the same dimensional space, we proceed to model the interactions between features and their contributions to the prediction target. The critical issue is how to fuse features from various sources to offer complementary information to each other. Given the recent significant performance of multi-head self-attention networks [vaswani2017attention] in modeling complex relationships, we utilize a multi-head attention mechanism to accomplish the interaction and fusion of features, employing different heads for different representation subspaces. First, we flatten the instance-wise feature embeddings () into a long-term relationship vector , then we feed it to the multi-head attention module for feature interaction and combine outputs concatenated from all heads (in the study, head number set as 2):
@@
To comprehensively evaluate the effectiveness of our proposed method, we selected two different classes of methods for comparative analysis based on the various representations of features: (i) Machine learning methods utilizing discrete features, for which we chose the most commonly used Xgboost [chen2016Xgboost]. (ii) Deep learning methods using continuous features, including classification networks such as the CNN architecture ResNet34 [he2016deep], the CNN-Attention architecture Tri2D-Net [chao2021deep], and the Transformer architecture ViT-B [dosovitskiy2020image]. Tri2D-Net is currently the leading method for CVD risk prediction on the LDCT-NLST dataset. The other category is based on multi-task learning approaches, with prior research [caruana1997multitask] demonstrating that models discover more general feature representations in solving multiple tasks, thereby improving generalization to unseen data. Additionally, we are aware that the nnUNet framework [isensee2021nnu] possesses strong data preprocessing and augmentation capabilities. Therefore, we integrated classification heads with fully connected layers and average pooling operations into segmentation networks such as nnUNet and nnFormer [zhou2023nnformer]. These heads receive input from the deepest feature layer of the encoder. Outputs provided by the multi-task methods include predicted cardiac segmentation masks and CVD risk classification probabilities. Except for Tri2D-Net, which directly uses an open-source model111https://github.com/DIAL-RPI/CVD-Risk-Estimator trained on the LDCT-NLST training set, all other comparison methods are trained from scratch based on the LDCT-NLST training set.

### v2

@@ 0.6 Deep continuous feature and discrete biomarker extraction
The extraction of deep continuous features relies on a pre-trained CVD risk prediction model, which has been trained on the LDCT-NLST training set. The model is based on a progressive coarse-to-fine framework. In the coarse stage (heart localization), a lightweight cardiac segmentation model is applied to localize the volume of the region of interest in the heart. The input chest LDCT scan is cropped into a smaller volume to reduce the influence of irrelevant content noise and conserve computational resources. Moving to the fine stage, the cropped volume is used to train a classifier based on ResNet34 [22], tasked with distinguishing between CVD-Positive and CVD-Negative cases. Ultimately, the outputs from the layer preceding the final fully connected layer of ResNet34 are extracted. These high-dimensional embeddings are rich in features relevant to CVDs, denoted as (, here ), and serve as continuous deep features of our proposed method.
@@
The CT quantitative biomarkers extraction process utilizes four specialized body part segmentation models, each trained on the 400 internal chest CT scans. These models are designed to segment different anatomical structures, including the heart chambers and pericardium, aortic and coronary calcium, aortic structure, and left and right lungs. The foundational architecture for these models is MedFormer [28]. These four segmentation models are applied to segment the whole LDCT-NLST dataset, with the results being reviewed by two radiologists and revised if necessary. The overall failure rate is less than 0.5. Fig. 4 visually showcases the results obtained from these fully automated body part segmentation models.
@@
We have established stable quantitative measures for each tissue composition based on the automated segmentation results without additional learning or adjustment. A total of quantitative biomarkers are calculated, with certain biomarkers such as the coronary artery calcium score (CACS) and the cardiothoracic ratio (CRT) having established associations with CVD and mortality in previous research [12, 27, 29, 30, 31]. Others are characterized by physicians using the results of the four segmentation models. All of these biomarkers are scalar and are symbolized as (where ).
@@
(i) Based on the segmentation results of the pericardium, three quantitative biomarkers related to pericardial fat are calculated: PFATV, PFATM, and PFATSTD. PFATV primarily serves to quantify the volume of pericardial fat. Initially, we used the result of pericardial segmentation (here marked in red in 4f) to identify the location of the pericardium on the CT scan. Subsequently, we further delineate the regions of pericardial fat (as depicted in 4b) within the Hounsfield Unit (HU) range of [-190HU, -30HU] [31]. Following this, we count the number of voxels within the 3D pericardial fat mask (illustrated in 4c), and by multiplying this count with the voxel volume ( resolution resolution resolution), we derive the value for PFATV, expressed in . Finally, we compute the mean and standard deviation of the attenuation intensity values for all voxels within the pericardial fat regions, denoted as PFATM and PFATSTD, respectively.
@@
(ii) Based on the segmentation results of calcifications (Coronary Artery Calcification and Thoracic Aorta Calcification), four calcification-related quantitative biomarkers are calculated: CACS, CACV, ACS, and ACV. CACS and ACS are primarily derived from an area perspective, calculated according to the Agatston Score [32] for coronary and thoracic aortic calcification scores, respectively. CACV and ACV, on the other hand, are calculated from a volumetric perspective, quantifying the volume of coronary calcification and thoracic aortic calcification, respectively. First, the calcification segmentation results are refined, with only regions with CT attenuation intensity greater than or equal to 130HU retained as the final calcification segmentation mask. Then, calculating calcification scores, the results are uniformly reconstructed to 3mm before computing CACS and ACS to minimize the impact of different CT slice thicknesses. In the calculation of calcification volume, the number of voxels within the 3D coronary calcification mask (green areas in Fig. 4d and Fig. 4e) and thoracic aorta calcification mask (red regions in Fig. 4d and Fig. 4e) are directly tallied, and then multiplied by the voxel volume ( resolution resolution resolution) to obtain CACV and ACV (in ).
@@
(iii) Based on the segmentation results of the thoracic aorta, three quantitative biomarkers characterizing the morphology of the thoracic aorta are computed: ATI, AMD, and AMDSTD. ATI represents the curvature of the thoracic aorta [29], while AMD and AMDSTD describe the overall maximum diameter of the aorta and its variation along the entire length. To calculate ATI, we first extract the centerline of the thoracic aorta using the segmentation outcome, and then determine the locations of the root point and end point (referred to as Point and Point in Fig. 4f). We then calculate ATI as the ratio of the length of the centerline (in ) to the straight linear distance between the point and the point (also in ). Subsequently, we generate cross-sectional views at 1 intervals along the centerline, with Fig. 4i and Fig. 4j exemplifying the cross-sectional views generated at positions and in Fig. 4(f). For each cross-sectional view, we measure the maximum diameter of the thoracic aorta, denoted as (where = centerline length/interval (1 ) + 1, it represents the total number of cross-sectional views). Finally, we compute AMD as the maximum value in (AMD = ) and ANDSTD as the standard deviation of the diameters in (AMDSTD = ).
@@
(iv) Based on the results of the heart segmentation (cardiac chambers and pericardium), four quantitative biomarkers are computed to characterize the morphology and structure of the heart: CHR, CLD, CSD, and CTR [33]. CHR is defined as the ratio of the volume of the cardiac chambers (the green regions in Fig. 4f) to the total volume of the entire heart (the combined red and green areas in Fig. 4f). Using the segmentation results of the heart, we select the maximal four-chambers axial view of the heart and perform an elliptical fitting on this segmented image to derive the cardiac long diameter (CLD) and short diameter (CSD). In Fig. 4f, half of line represents the CSD while half of line represents the CLD. Based on the calculation method described in the previous study [29], we determine the width of the heart, represented by line in Fig. 4g. Similarly, utilizing the segmentation results of the lungs (as seen in Fig. 4l), we obtain the width of the entire lung, denoted by line in Fig. 4g. Finally, CTR is calculated as the ratio of line to line .
@@
Recent studies have shown that methods that integrate features from various sources into one representation can offer complementary information to each other and lead to better performance [24]. However, not all features contribute to the prediction of the target, and certain artificially designed image quantitative biomarkers may actually hinder performance. Furthermore, the relevance and specific contributions of the features to the output target are typically unknown. Therefore, we have developed a features joint representation module (Fig. 5a) which can help identify which features are most significant for the prediction problem, reinforcing the most relevant features for CVD risk prediction, and suppressing any unnecessary features that could have a negative impact on performance. This module comprises an instance-wise feature-gated mechanism and a soft instance-wise feature interaction mechanism.
@@
To improve the expressive power of the model and better capture the relationship between features and the output target, we apply non-linear processing to each instance-wise feature embedding . However, determining the extent of required non-linear processing remains a complex task. Some studies [34, 35] suggest that simpler models may benefit in datasets with noise. Given that our discrete quantitative biomarkers are automatically derived from the four pre-trained body part segmentation models without any additional adjustment, it implies that the input contains unavoidable noise. In consideration of this, we employ the GRN proposed in [19], which is a notably flexible and simple architecture (Fig. 5b), as the non-linear operation.
Where is the fully-connected Layer, is the Exponential Linear Unit activation function [36], is the Gated Linear Units [37] and is the standard layer normalization [38]. During training, dropout is applied, and the dropout rate is set as 0.5. For each instance-wise feature embedding , a non-linear operation is employed by its own GRN:
@@
Once the continuous and discrete features are aligned into the same dimensional space, we proceed to model the interactions between features and their contributions to the prediction target. The critical issue is how to fuse features from various sources to offer complementary information to each other. Given the recent significant performance of multi-head self-attention networks [20] in modeling complex relationships, we utilize a multi-head attention mechanism to accomplish the interaction and fusion of features, employing different heads for different representation subspaces. First, we flatten the instance-wise feature embeddings () into a long-term relationship vector , then we feed it to the multi-head attention module for feature interaction and combine outputs concatenated from all heads (in the study, head number set as 2):
@@
To comprehensively evaluate the effectiveness of our proposed method, we selected two different classes of methods for comparative analysis based on the various representations of features: (i) Machine learning methods utilizing discrete features, for which we chose the most commonly used Xgboost [21]. (ii) Deep learning methods using continuous features, including classification networks such as the CNN architecture ResNet34 [22], the CNN-Attention architecture Tri2D-Net [18], and the Transformer architecture ViT-B [39]. Tri2D-Net is currently the leading method for CVD risk prediction on the LDCT-NLST dataset. The other category is based on multi-task learning approaches, with prior research [40] demonstrating that models discover more general feature representations in solving multiple tasks, thereby improving generalization to unseen data. Additionally, we are aware that the nnUNet framework [41] possesses strong data preprocessing and augmentation capabilities. Therefore, we integrated classification heads with fully connected layers and average pooling operations into segmentation networks such as nnUNet and nnFormer [42]. These heads receive input from the deepest feature layer of the encoder. Outputs provided by the multi-task methods include predicted cardiac segmentation masks and CVD risk classification probabilities. Except for Tri2D-Net, which directly uses an open-source model111https://github.com/DIAL-RPI/CVD-Risk-Estimator trained on the LDCT-NLST training set, all other comparison methods are trained from scratch based on the LDCT-NLST training set.